OpenStudy (anonymous):

The quadratic functiob with vertex (1,3) is shown below find the domain and range

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

the domain of a quadratic function is all real numbers

OpenStudy (anonymous):

They only give a graph and the vertex

OpenStudy (misty1212):

without the picture it is not clear what the second answer is, depends on whether it goes up or down

OpenStudy (anonymous):

It goes down

OpenStudy (anonymous):

Passes through points 0,1 and 2,1

OpenStudy (misty1212):



OpenStudy (anonymous):

How do you do that

OpenStudy (misty1212):

draw tool belowo

OpenStudy (misty1212):

is that what it looks like?

OpenStudy (anonymous):

Yea

OpenStudy (misty1212):

then the largest y can be is 3, can't be any bigger

OpenStudy (misty1212):

so range is \((-\infty, 3]\) or \(y\leq 3\) depending on how you are supposed to write it

OpenStudy (anonymous):

What about domain

OpenStudy (misty1212):

i think i said that already right?

OpenStudy (anonymous):

They didnt give me a quadratic function

OpenStudy (misty1212):

unless you impose some sort of restriction , the domain of any quadratic function is all real numbers \(\mathbb{R}\) or \((-\infty, \infty)\)

OpenStudy (misty1212):

you only need the graph to find the highest point on it